ACCOUNT NAME DR CR CASH 38 9 ACCOUNTS RECEIVABLE PREPAID INSURANCE EQUIPMENT 80 ACCUMULATED DEPRECIATION ACCOUNTS PAYABLE 9 C: What is Net Income for Mint Cleaners?What is ending Retained Earnings for Mint Cleaners?What are total assets for Mint Cleaners?
Answer #1

1) calculation of net income for mont cleaners:

Net income = sales revenue - supplies expenses

= 80 - 26

Net income = 54

2) calculation of ending retained earnings:

Ending retained earnings

= opening retained earnings + net income - dividend paid

= 4 + 54 - 0

Ending retained earnings = 58

3) calculation of total assets for mint cleaners:

Total Assets = Cash + Accounts receivable + Equipment

= 38 + 9 + 80

=127.